Soybean meal is 18% protien, cornmeal is 9% protien. How many pounds of each should be mixed together in order to get 360 lb mixture that is 13% protien?
How many pounds of cornmeal should be in the mixture?
How many pounds of soybean meal should be in the mixture?

Answer by checkley77(12844)

.18x+.09(360-x)=.13*360
.18x+32.4-.09x=46.8
.09x=46.8-32.4
.09x=14.4
x=14.4/.09
x=160 lb. of soybeans is used.
360-260=200 lb. of cornmeal is used.
Proof:
.18*180+.09*300=46.8
28.8+18=46.8
46.8=46.8